A person needs $5,000 in four years. The person can afford to deposit $275 every three months into an account. Determine to the nearest .01% the nominal rate of interest converted quarterly that is needed to accumulate $5,000.
Answer #1

Ans 6.72%

P = Periodic payments
r = rate of interest
n = no of years
Future Value of Annuity = P ( (1 + r)n - 1 ) / r
5000 = 275* ((1 + r%/4)^16 - 1) / (r%/4)
5000 / 275 = ((1 + r%/4)^16 - 1) / (r%/4)
18.1818181818182 = ((1 + r%/4)^16 - 1) / (r%/4)
r = 6.72%